a college paid $1748 for 46 calculators. They paid $9 for each scientific calculator. all other graphing calculators cost $55 each how many of each calculator was ordered
Answer by nerdybill(7384)

a college paid $1748 for 46 calculators. They paid $9 for each scientific calculator. all other graphing calculators cost $55 each how many of each calculator was ordered
.
Let x = number of scientific calculators
then
46-x = number of graphing calculators
.
9x + 55(46-x) = 1748
9x + 2530 - 55x = 1748
2530 - 46x = 1748
-46x = -782
x = 17 (number of scientific calculators)
.
graphing calculators:
46-x = 46-17 = 29